OpenStudy (anonymous):

Elizabeth has worked for 5 months and earned 6 vacation days. If she continues to earn vacation days at this rate, how many more months does she need to work to reach a total of 10 vacation days?

OpenStudy (anonymous):

i have to do that for my school on a website called fcat focus and i looked up the question and found this-_-

OpenStudy (anonymous):

5/6 = ?/10 10*5= 50 50/6= 8.3